From the given table calculate TR, MR and AR and highlight the relationship between

total revenue (TR) and marginal revenue (MR). (10 Marks)

Quantity Price

50 200

60 150

70 100

80 50

90 10


1
Expert's answer
2021-06-29T10:05:02-0400

Total revenue (TR) = Quantity X price

50x200 = 10000

60 x150 = 9000

70 x100 = 7000

80 x50 = 4000

90 x10 = 900

Marginal Revenue = Change in Total Revenue / Change in Quantity Sold

                                       = (10000-9000)/ (60-50)

                                       = 1000/10

                                       = 100

                               

                                 = (9000-7000)/ (70-60)             

                                 = 2000/10

                                 = 200

 

                                 = (7000-4000)/ (80-70)

                                 = 3000/10

                                 = 300

                                 

                                 = (4000-900)/ (90-80)

                                 = 3100/10

                                 = 310

 

Average revenue = Total revenue / quantity of units

                               =200/1

                                = 200

                                = 150/ 1

                                 = 150

                                 = 100/1

                                 = 100

                                         = 50/1

                                          = 50

                                           = 10/1

                                            = 10

Relationship between Total Revenue (TR) and Marginal Revenue (MR)

When Price Falls with rise in output, firm follows its own pricing policy. However, it can increase sales only by reducing the price